OpenStudy (anonymous):

The midpoint of a line segment with end points as (-10, y1) and (-6, 7) is (-8, 6). What is the value of y1? Midpoint formula is (x1 + x2)/2 , (y1 + y2)/2 but I keep getting a wrong answer.

OpenStudy (amistre64):

write down the number 6, thats and end write down the number 7, thats halfway to where youwant to be whats the next number that would get you all the way?

OpenStudy (anonymous):

(y + 7) /2 = 6 I solve for y, right^

OpenStudy (amistre64):

no

OpenStudy (amistre64):

the midpoint formula uses ends to find middle; you got a middle and end mixed up

OpenStudy (amistre64):

6, 7, y , lets move this to the start at 0 -6 -6 -6 -------- 0 , 1 , n ; since you moved 1 to get halfway, move 1 again to get to the end; n = 1+1 = 2 , now move this all back to start at 6 :) 0 , 1 , 2 +6 +6 +6 ----------- 6 , 7 , 8

OpenStudy (amistre64):

its not the "formula", but it is the concept behind finding the middle and extremes

OpenStudy (phi):

(y + 7) /2 = 6 I solve for y, right^ yes, this is one way to find the answer another way is notice that if you start at one end (-6,7) and move (-2,-1) you get to the mid point (-8,6). move the same amount (-2,-1) from the mid point to get to the other end.

OpenStudy (anonymous):

Oh okay, now I understand. Thank you! (:

OpenStudy (anonymous):

Sorry I took so long to check back!

OpenStudy (anonymous):

So y1 would 5, correct?

OpenStudy (amistre64):

... i seem to have gotten the points all jumbled up when reading the question :) yes. 5 is it

OpenStudy (anonymous):

Thanks. (: